$SPAD/src/input folium.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(curve((t**2-1)/(3*t**2+1),t*(t**2-1)/(3*t**2+1)),_
     t = -3..3, [title "Folium of Descartes"])

———-

References^{[1] nothing}

4